OpenStudy (anonymous):

What is the solution to the equation 2 1 + 2x = 3 ? x ≈ –0.292 x ≈ –0.185 x ≈ 0.185 x ≈ 0.292

OpenStudy (anonymous):

2^1 + 2x = 3

OpenStudy (anonymous):

2x=3-2 x=1/2

OpenStudy (anonymous):

but this answer is not in ur choices... whats going on?

OpenStudy (anonymous):

its 2^ 1 + 2x = 3 I forgot to put the symbol

OpenStudy (anonymous):

2^(1+2x)=3 or (2^1)+2x=3

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

oh ok. then use log base 2 on both side. (1+2x)=lob(base2)3 x=(log(base2)3)-1)/2 x=.2924815